OpenStudy (anonymous):

Find the length of the missing side of the right triangle. (1 point) 25 ft 31 ft 22.96 ft 625 ft

jhonyy9 (jhonyy9):

what are the side length what you know ?

OpenStudy (vhtran):

\[a^2+b^2=c^2\]

OpenStudy (vhtran):

have you learned about that equation before?

OpenStudy (anonymous):

yes

OpenStudy (vhtran):

c should be the Hypotenuse \[24^2+7^2=c^2\]

OpenStudy (anonymous):

Okay

OpenStudy (vhtran):

576+49=c^2

OpenStudy (anonymous):

625?

OpenStudy (vhtran):

\[625=c^2\] You are trying to find c, square root both sides to get c.

OpenStudy (vhtran):

25=c

OpenStudy (anonymous):

Thats what i got thankss

OpenStudy (vhtran):

np